Circular No. 7701

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 1996cl, 1996cp, AND 1996cq Further to IAUC 7700, Gal-Yam et al. write: “A comparison of HST

images of the galaxy cluster MS 1054.4–0321 (z = 0.83) obtained during 1996 Mar. 13 (total exposure 18 200 s through the F814W filter) with later images obtained during 1998 May 29 (total exposure 6600 s, same filter) reveals two new sources. SN 1996cp at α = 10^{h}57^{m}03^{s}.5, δ = −3^{o}37′41′′ (equinox 2000.0) is detected at magnitude [I814] = 26.5 (equivalent to Cousins I = 26.2) and is located at the same declination and 0′′.9 east of an unidentified host galaxy. SN 1996cq (at α = 10^{h}57^{m}02^{s}.2, δ = −3^{o}37′34′′) is detected at [I814] = 26.0 (equivalent to Cousins I = 25.7). A very faint fuzz near the image detection limit may be associated with the host galaxy of this object. Both sources are absent from the 1998 data (limiting I814 mag at least 27.7). Both new objects have stellar PSFs at the HST resolution limit and appear on six individual sub-exposures obtained over a period of 10.5 hr. We estimate an upper limit of 0′′.12/day on any motion of these sources. We note that the comparison of the 1996 images of this cluster with the 1998 data also clearly shows SN 1996cl (IAUC 6621), 0′′.05 west and 0′′.1 south of its host galaxy’s nucleus; SN 1996cl appears at [I814] = 23.8 (equivalent to Johnson-Cousins I = 23.5), consistent with the magnitude reported by Perlmutter et al. (1998, Ap.J. 517, 565).”

SUPERNOVA 2001bf IN MCG +04-42-22 R. Chornock, A. V. Filippenko, and W. D. Li, University of Califor-

nia at Berkeley, write: “Inspection of CCD spectra (range 330–1000 nm), obtained on July 27 UT with the Shane 3-m reflector at Lick Observatory, reveals that SN 2001bf (IAUC 7620) is of type Ia (as suggested on IAUC 7625), not Ic (IAUC 7636). Our classification is supported by the spectrum

obtained by the CfA group (http://cfa-www.harvard.edu/cfa/oir/
Research/supernova/spectra/).”

SUPERNOVA 2001dw IN NGC 1168 Corrigendum. On IAUC 7691, line 6, for 13′′ west and 9′′.5 south read

13′′ east and 9′′.5 north

COMET 19P/BORRELLY Visual m_{1} estimates by N. Biver: Aug. 12.12 UT, 11.1 (Ablis, France,

0.26-m refl.); 18.13, 10.7; 26.13, 10.4 (Noordwijk, Netherlands, 0.20-m refl.).

2001 August 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green